### v1

# Optimal sequencing depth for single-cell RNA-sequencing in Wasserstein space

Abstract.
How many samples should one collect for an empirical distribution to be as close as possible to the true population? This question is not trivial in the context of single-cell RNA-sequencing. With limited sequencing depth, profiling more cells comes at the cost of fewer reads per cell. Therefore, one must strike a balance between the number of cells sampled and the accuracy of each measured gene expression profile. In this paper, we analyze an empirical distribution of cells and obtain upper and lower bounds on the Wasserstein distance to the true population. Our analysis holds for general, non-parametric distributions of cells, and is validated by simulation experiments on a real single-cell dataset.
Key words and phrases:
RNA sequencing, optimal transport, Wasserstein distance, optimal depth, empirical distribution, nonparametric inference
2020 Mathematics Subject Classification:
62G05, 49Q22, 62D99
1. Introduction
The recent success of single-cell RNA sequencing (scRNA-seq) technologies [kleinDropletBarcodingSingleCell2015, macoskoHighlyParallelGenomewide2015] has resulted in massive, high-dimensional datasets, which invite interesting statistical questions. There has been intensive work on clustering of cells [traagLouvainLeidenGuaranteeing2019, grabskiSignificanceAnalysisClustering2023], transformations and variance stabilization [ahlmann-eltze_ComparisonTransformationsSinglecell_2023], trajectory inference [schiebingerOptimalTransportAnalysisSingleCell2019, qiuSystematicReconstructionCellular2022, lavenant_MathematicalTheoryTrajectory_2023] and manifold learning and denoising [vandijkRecoveringGeneInteractions2018, huangSAVERGeneExpression2018]. Mathematically, the basic input in each of these tasks is an empirical distribution of cells; these are the cells sampled by scRNA-seq. However, contrary to the ordinary setting of high-dimensional statistics, the support of the empirical distribution is also noisy. This is because cells’ gene expression profiles are measured, through sequencing, with finitely many ‘reads’.
A ‘read’ is the fundamental unit of information in DNA sequencing, and in RNA sequencing for that matter. The relative abundances of gene transcripts, which characterize cell type, can be measured by extracting RNA from the cell, and reading the sequence of nucleotide letters (ACTG) using a DNA sequencing machine. Each ‘read’ documents the presence of one molecule of RNA in one cell. Early forms of single-cell RNA sequencing produced high-fidelity measurements of gene expression by profiling small numbers of cells, with a large number of reads per cell. Droplet-based technologies made it possible to profile vastly more cells, but with fewer reads per cell. It soon became clear that this was beneficial for identifying rare cell types in complex populations of cells [jaitinMassivelyParallelSingleCell2014, shalekSinglecellRNAseqReveals2014, pollenLowcoverageSinglecellMRNA2014, streetsHowDeepEnough2014, bacher2016design, haque2017practical, ecker2017brain, dal2019design].
Here, we analyze the natural experimental design question: given a finite budget of reads, how many cells should be profiled so that the empirical distribution is as close as possible to the true population? As illustrated in Figure 1, profiling fewer cells allows each expression profile to be measured more accurately (with more reads per cell), but the full population is not well explored. On the other hand, profiling more cells gives an empirical distribution with larger support, but with more noise in the support which can blur out the finer features of the population distribution such as the two clusters (the peaks in the shaded contour plots).
[…]
Figure 1. The trade-off between deep and shallow sequencing. In this schematic, we have three panels illustrating the effect of different sequencing depths. In each panel, the shaded contours describe the unknown distribution of gene expression in a population of cells. The green dots represent the true gene expressions of sampled cells, whereas the red markers indicate the corresponding measured gene expressions obtained via the sequencing process. The yellow lines match a cell’s true and measured gene expressions. When the number of cells is very low (left panel), each cell receives a high proportion of the budget of reads, leading to low measurement noise. However, the population distribution is not well-captured. Conversely, when the number of cells is very high (right panel), each cell receives a small fraction of the budget, resulting in very noisy measurements. The central panel depicts an intermediate case that balances these two effects.
In this work, we analyze this experimental design question in a very general setting. We leverage optimal transport, more precisely Wasserstein distance, to evaluate the discrepancy between between the noisy empirical distribution obtained by RNA sequencing and the unknown ground truth. There are many advantages of using optimal transport. First, it is well defined for the comparison between sufficiently general probability distributions. Moreover, Wasserstein distance takes into account not vertical but horizontal difference between distributions, which displays geometric information of distributions. Second, Wasserstein geometry naturally involves gradient flow, the evolution of distributions [mccann1994convexity, McCann1997Adv, Benamou_brener2000, otto2001geometry]. Since the recent demands of trajectory inference in single-cell RNA-sequencing data have emerged [schiebingerOptimalTransportAnalysisSingleCell2019, qiuSystematicReconstructionCellular2022, lavenant_MathematicalTheoryTrajectory_2023], Wasserstein geometry and Wasserstein gradient flows are the state-of-art mathematical tool for this field. Lastly, Wasserstein distance is computationally tractable. In his seminar work, Cuturi [cuturi2013sinkhorn] proposed an almost linear-complexity approximation algorithm, the so-called Sinkhorn algorithm or entropic optimal transport, which allows practitioners to use it in reality.
In this flexible framework, we are able to obtain results for general distributions of cells, without any assumptions on the distribution of interest. Moreover, because we focus on the empirical distribution directly, and not on any specific estimator, we hope that our results can inform broad estimation tasks. Our result depends on the sparsity of gene expression and the intrinsic low dimensionality of its distribution explicitly, which explains the role of those factors in RNA sequencing in a clear way. Furthermore, we conduct simulation experiments on a real single-cell dataset to validate and support the theory.
1.1. Setup and contributions
In this section, we introduce the mathematical notation and the setting of the problem, and state our contribution. Throughout this paper, $d$, $n$ and $m$ denote the number of genes, i.e. the dimension of gene expression vector, the number of cells, and the number of reads, respectively. In the sequel we use big-$O$ notation as usual, $f(n)\sim g(n)$ if $f$ and $g$ have the same first order growth, and $f(n)\lesssim g(n)$ if $f(n)$ is less than $g(n)$ in terms of variables of interest while other universal constants are fixed.
We model the measurement process with two steps. The first step is the selection of $n$ cells from the population. This will give us an empirical distribution of cells $\mu_{n}$, defined precisely in (1.1) below. Then, in the second step, we measure the gene expression of each cell through sequencing to obtain a noisy empirical distribution $\widehat{\mu}_{n}(m)$, defined precisely in (1.2) below.
In sequencing, only the relative abundance of gene products matters. Therefore, we represent gene expression profiles of cells in relative terms. We denote a (relative) gene expression vector $P=(p_{1},\dots,p_{d})\in\Delta_{(d-1)}$ where $\Delta_{(d-1)}$ is the $(d-1)$-dimensional simplex. Since each cell is represented by its gene expression; hence it can be regarded as $P\in\Delta_{(d-1)}$. Let $\mu$ be the unknown true distribution of gene expression $P$, i.e., $P$ is distributed according to $\mu$. Our goal is to obtain a distribution which approximates $\mu$.
In the first step of the measurement process, we sample $n$ cells from the population. This gives us i.i.d. samples from $\mu$ which we denote by $\boldsymbol{P}:=\{P_{1},\dots,P_{n}\}\subseteq\Delta_{(d-1)}$. We denote the unknown true empirical distribution by
$\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{P_{i}}.$ (1.1)
In the second step of the measurement process, we consider two related models for the sequencing procedure. In both cases, cell expression profiles are generated multinomially from the true expression profile. The two models differ in how many reads are selected per cell. In the first model, cell sampling frequency $U\in\mathbb{R}_{>0}$ can depend on gene expression profile $P$. In the second model, $P$ and $U$ are independent. We use $\mu_{(P,U)}$ to denote the joint distribution of $U$ and $P$ for the fist model. When we use $\mu$ only, it always denotes the (marginal) distribution of $P$, not of $U$.
Let $\{U_{1},\dots,U_{n}\}$ be the sampling frequencies of the $n$ cells sampled in step 1. We set $u_{i}:=\frac{U_{i}}{\sum_{j}U_{j}}$ for $i=1,\ldots,n$, and refer to $\{u_{1},\dots,u_{n}\}$ as (normalized) cell weights. One can regard $\boldsymbol{u}:=(u_{1},\dots,u_{n})\in\Delta_{(n-1)}$ as a probability measure over the $n$-cells selected in step 1 which will now be sequenced in step 2. It specifies how often a read is selected for each cell: for each iteration $I$-th cell is chosen identically and independently according to $\boldsymbol{u}$. Given $m$ reads and $n$ cells, suppose $I$-th cell is chosen independently and randomly according to $\boldsymbol{u}$ at $t$-th iteration. Let $\{e_{1},\dots,e_{d}\}$ be the standard basis of $\mathbb{R}^{d}$. A coordinate of gene expression, $e_{j}$, is counted independently and randomly according to $P_{I}=(p_{I1},\dots,p_{Id})$, and recorded by using one read. After $m$ iterations (since we have only $m$ reads), the total counts are summed and normalized. One then obtains the collection of measured (relative) gene expressions of $n$ cells by $m$ reads $\{\widehat{P}_{1},\dots,\widehat{P}_{n}\}\subseteq\Delta_{(d-1)}$, and the noisy empirical distribution is obtained as
$\widehat{\mu}_{n}(m):=\frac{1}{n}\sum_{i=1}^{n}\delta_{\widehat{P}_{i}}.$ (1.2)
Algorithm 2 describes shallow sequencing formally. Note that variants of such a multinomial sequencing model are standard in the literature [townesFeatureSelectionDimension2019, heimbergLowDimensionalityGene2016], and are closely related to popular Poisson model [wangGeneExpressionDistribution2018, zhang2020determining].
Then, the goal is to answer two questions on $\widehat{\mu}_{n}(m)$, the output of shallow sequencing, in order to promise the closeness of $\widehat{\mu}_{n}(m)$ to $\mu$:
  1. (1)
    given $n$ cells, how many reads at least should we use to obtain good $\widehat{\mu}_{n}(m)$,
  2. (2)
    given $m$ reads, how many cells at most can we use to obtain the best $\widehat{\mu}_{n}(m)$.
In Theorem 2.4 and Theorem 2.5, we resolve question (1) by showing that $m\gtrsim\frac{n}{\varepsilon^{2}}$ is sufficient to guarantee that Wasserstein distance between $\widehat{\mu}_{n}(m)$ and $\mu_{n}$ is less than $\varepsilon$, and the expected Wasserstein distance between $\widehat{\mu}_{n}(m)$ and $\mu_{n}$ is lower bounded by $\frac{n}{m}$ in general. Built from these results, we show that $n\sim m^{1-o_{k}(1)}$ is optimal in Theorem 2.6, Corollary 2.7 and Corollary 2.8, which solves question (2). In particular, our results depend on the sparsity and the intrinsic dimension explicitly, which illustrates their effectiveness to approximation.
1.2. Comparison to related work
[…]
$\frac{m}{n}\sim\frac{\kappa^{2}}{n\lambda_{i}\varepsilon^{2}}$ (1.3)
where $\varepsilon>0$ is the desired error such that $\|v_{i}-\widehat{v}_{i}\|_{2}\leq\varepsilon$, and $\kappa$ is a constant that can be estimated from existing data. The allocation (1.3) is interesting since the number of cells $n$ cancels out, having no direct effect on the eigenvector error. Nevertheless, (1.3) does not address our problem of estimating the population distribution $\mu$ or its associated quantities, since $v_{1},\dots,v_{d}$ are the principal directions of the empirical distribution $\mu_{n}$ and not of the population $\mu$. In that direction, it may be more meaningful to compare $\widehat{\Sigma}_{n}$ with the population covariance $\Sigma$ (i.e. covariance of $\mu$).
[…]
$\frac{m}{nd}=\Theta(1)$ (1.4)
(see [zhang2020determining, Theorems 1 and 2]), which we rewrite as the linear relationship $n=\Theta(m)$ by hiding the constant $d$ for simplicity. This is slightly different from our allocation, which is mildly sublinear. Indeed, their work addresses a slightly different problem: parametric estimation of various statistics of $\mu$, and additionally under an exponential family assumption, estimation of the 1-dimensional marginals of $\mu$. This is further manifested in the estimation rates, expressed in terms of the budget $m$. Given a 1-dimensional marginal of $\mu$ parametrized by some $\theta\in\mathbb{R}^{p}$, they establish the parametric rate
$\mathbb{E}\|\widehat{\theta}^{\mathrm{EB}}-\theta\|_{2}^{2}=O\left(\frac{1}{m}\right)$
under (1.4), where $\widehat{\theta}^{\mathrm{EB}}$ is the empirical Bayes estimator of $\theta$.
We study the optimal allocation problem in a fully non-parametric setting, for the first time to our knowledge. Our allocation $n\sim\left(\frac{m}{\mathbb{E}\Arrowvert P\Arrowvert_{0}}\right)^{1-\frac{2}{k+2}}$ (see Corollary 2.7) yields the rate
$\mathbb{E}W_{p}(\widehat{\mu}_{n}(m),\mu)\leq O\left(\frac{\mathbb{E}\left\|{P}\right\|_{0}}{m}\right)^{\frac{1}{2+k}}$
[…]
1.3. Organization
[…]
2. Main results
2.1. Preliminaries
In the present paper, Wasserstein distance is the quantity to measure the goodness of the noisy empirical distribution.
Let’s discuss the history of optimal transport very briefly. Optimal transport(OT) has a long story and is an interdisciplinary field connecting to both pure mathematics and applied sciences. Proposed by Monge [monge1781memoire], optimal transport was developed by Kantorovich [kantorovich1942translocation, kantorovich1948problem] developed a standard framework for analyzing optimal transport, for which he received Nobel prize in economic sciences. Since then, OT has related to PDE, geometry and probability [Brenier91, mccann1994convexity, Caffarelli_contraction, Benamou_brener2000, Otto2000GeneralizationOA, otto2001geometry, Lott2004RicciCF, Figalli2010Invent, Kim2010JEMS, Figall2011duke, Caffarelli2013Reine, Figalli2015CMP, GOZLAN20173327]. Recently, OT has also played a central role in finance, statistics and machine learning: model free approach in mathematical finance and its statistical inference [hobson2012robust, Beiglb_Juillet16, kim2024convexorder], optimization of mean field two-layer neural networks [mei2018mean, chizat2018global, sirignano2020mean, Rotskoff2022CPAM], diffusion model [de2021diffusion, wang2021deep, hamdouche2023generative], variational inference [lambert2022variational, diao2023forward] and adversarial training [pydi2021the, Trillos2020AdversarialCN, jakwang2023JMLR].
Given spaces $\mathcal{X}$ and $\mathcal{Y}$, and “cost function” $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}\cup\{-\infty,+\infty\}$, the Monge problem (named for a special case proposed in [monge1781memoire]) is to find a map $T:\mathcal{X}\to\mathcal{Y}$ which achieves the minimum of the following optimization problem:
[…]
where $(T)_{\#}(\mu)$ is a pushforward measure, i.e. $(T)_{\#}(\mu)(B):=\mu\left(T^{-1}(B)\right)$ for any measurable function $T$ and measurable subset $B$. It is well known that, however, such a $T$ may not exist even for nice $c$. For the well-definedness of the problem, Kantorovich extended the solution space of the Monge problem to the set of couplings of marginals [kantorovich1942translocation, kantorovich1948problem]. The problem becomes an infinite dimensional linear optimization problem, which is defined as
[…]
If the spaces $\mathcal{X}$ and $\mathcal{Y}$ are Polish(complete and metrizable), then $\Pi(\mu,\nu)$ is non-empty and compact with respect to weak topology. For nonnegative and lower semi-continuous $c$, a minimizer of Kantorovich’s problem always exists [Oldandnew, Theorem 4.1].
The Kantorovich formulation provides a rich mathematical framework: it can be leveraged to define a distinguished family of metrics on the space of probability measures. Let $\mathcal{X}=\mathcal{Y}$ be a Polish (complete metric) space. The $p$-Wasserstein distance, or Kantorovich-Rubinstein distance, is defined over $\mathcal{P}_{p}(\mathcal{X})$, as follows:
${W}_{p}(\mu,\nu):=\min_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{X}\times\mathcal{X}}\mathrm{dist}(x,y)^{p}d\pi(x,y)\right)^{\frac{1}{p}}.$
[…]
Definition 2.1.
[Niles-Weed_Bach] Let $(\mathcal{F},||\cdot||)$ be a (semi)normed space. Given a set $S\subseteq\mathcal{F}$, the $\varepsilon$-covering number of $S$, denoted by $\mathcal{N}(\varepsilon,S,||\cdot||)$, is the minimum $m$ such that there exist $m$ closed balls $B_{1},\dots,B_{m}$ of diameter $\varepsilon$ such that $S\subseteq\cup_{i=1}^{m}B_{i}$. The centers of the balls need not belong to $S$. The $\varepsilon$-dimension of $S$ is defined as
$d(\varepsilon;S):=\frac{\log\mathcal{N}(\varepsilon,S,||\cdot||)}{-\log\varepsilon}.$
Given a probability measure $\mu$ on a metric space $\mathcal{F}$, the $(\varepsilon,\tau)$-covering number is
$\mathcal{N}(\varepsilon,\tau;\mu):=\inf\{\mathcal{N}(\varepsilon,S,||\cdot||):\mu(S)\geq 1-\tau\}$
and the $(\varepsilon,\tau)$-dimension is
$d(\varepsilon,\tau;\mu):=\frac{\log\mathcal{N}(\varepsilon,\tau;\mu)}{-\log\varepsilon}.$
The upper and lower Wasserstein dimensions are respectively,
$d^{*}_{p}(\mu)$ $:=\inf\left\{s\in(2p,\infty):\limsup_{\varepsilon\to 0}d(\varepsilon,\varepsilon^{\frac{sp}{s-2p}};\mu)\leq s\right\},$
$d_{*}(\mu)$ $:=\lim_{\tau\to 0}\liminf_{\varepsilon\to 0}d_{\varepsilon,\tau}(\mu).$
The main technique that we will leverage is to construct a simpler random variable which stochastically dominates a random variable of interest to us. Stochastic dominance is frequently used in various fields including finance, economics, statistics and etc. in order to measure uncertainty among several random variables(vectors). Among many variants of it, here we rely on the first order stochastic dominance.
Definition 2.2.
[roch2024modern, Definition 4.2.1] Let $\mu$ and $\nu$ be probability measures on $\mathbb{R}$. The measure $\mu$ is said to stochastically dominate $\nu$ if for all $x\in\mathbb{R}$,
$\mu\left((x,\infty)\right)\geq\nu\left((x,\infty)\right).$
A real random variable $X$ stochastically dominates $Y$ if the law of $X$ dominates the law of $Y$.
Let’s introduce the usual multinomial sampling scheme from which RNA sequencing is defined. Let $\{e_{1},\dots,e_{K}\}$ denote the standard basis of $\mathbb{R}^{K}$ and $Q=(q_{1},\dots,q_{K})\in\Delta_{(K-1)}$ be a probability distribution over $\{e_{1},\dots,e_{K}\}$, i.e. if $E$ is drawn by $Q$, then
$\mathbb{P}_{Q}\left(E=e_{k}\right)=q_{k}.$
The multinomial sampling proceeds as follows. At each iteration we draw $E\sim\{e_{1},\dots,e_{K}\}$ according to $Q$ identically and independently. After $T$ iterations, it outputs $\widetilde{Q}=\widetilde{Q}(T)$, which is the empirical distribution of $Q$. The below is the pseudocode of it.
$T\in\mathbb{N},Q\in\Delta_{(K-1)}$
Set $Z(0)=\boldsymbol{0}\in\mathbb{R}^{K}$
for $t=1,\dots,T$ do
     $E(t)\sim\{e_{1},\dots,e_{K}\}$ according to $Q$ independently
     Update $Z(t)=Z(t-1)+E(t)$
end for
Set $\widetilde{Q}:=\frac{Z(T)}{T}\in\Delta_{(K-1)}$.
Algorithm 1 Multinomial$(T,Q)$
We end this subsection to introduce $\ell_{0}$ norm, which is neither norm nor even pseudo-norm but is called norm conventionally. $\ell_{0}$ norm of $P$ is a critical parameter of the main results. For $P=(p_{1},\dots,p_{d})\in\mathbb{R}^{d}$, the $\ell_{0}$ norm of $P$ is defined as
$\Arrowvert P\Arrowvert_{0}:=\sum_{j=1}^{d}\mathds{1}_{p_{j}\neq 0}.$ (2.1)
One can understand $\ell_{0}$ norm as the sparsity of a vector.
2.2. Theorems
Let $\Delta_{(d-1)}$, the $(d-1)$-dimensional simplex, be equipped with $\ell_{q}$ distance for any $1\leq q\leq\infty$: for $P,P^{\prime}\in\Delta_{(d-1)}$
$\mathrm{dist}(P,P^{\prime}):=\Arrowvert P-P^{\prime}\Arrowvert_{q}=\begin{cases}\left(\sum_{j=1}^{d}|p_{j}-p^{\prime}_{j}|^{q}\right)^{\frac{1}{q}}&\text{ if $1\leq q<\infty$,}\\ \max_{1\leq j\leq d}|p_{j}-p^{\prime}_{j}|&\text{ if $q=\infty$.}\end{cases}$
Recall $\boldsymbol{P}=\{P_{1},\dots,P_{n}\}\subseteq\Delta_{(d-1)}$ and $\boldsymbol{u}=(u_{1},\dots,u_{n})\in\Delta_{(n-1)}$ denote the set of i.i.d. cells (samples, or gene expressions) and the weight vector over $\boldsymbol{P}$, respectively. At each iteration, $I$-th cell is chosen according to $\boldsymbol{u}$ independently, and $e_{j}$ is chosen according to $P_{I}$ independently and recorded. After $m$ iterations, all records are summed and normalized. In this sense, the shallow sequencing is a variant of Algorithm 1, which runs two-stage multinomial sampling. Notice that some of $P_{i}$’s never appear during the sequencing with positive probability. For such $P_{i}$’s, one just defines their approximations arbitrarily (though it is far from the best). Finally, the noisy empirical distribution, $\widehat{\mu}_{n}(m)$, is obtained. Its pseudocode is written below.
$m\in\mathbb{N}$, $\boldsymbol{P}\subseteq\Delta_{(d-1)}$, $\boldsymbol{u}\in\Delta_{(n-1)}$
Set $Z_{i}(0):=\boldsymbol{0}\in\mathbb{R}^{d}$ for $1\leq i\leq n$
for $t=1,\dots,m$ do
     Draw $I(t)\sim\{1,\dots,n\}$ according to $\boldsymbol{u}$
     $E(t)\sim\{e_{1},\dots,e_{d}\}$ according to $P_{I(t)}$
     Update $Z_{I(t)}(t)=Z_{I(t)}(t-1)+E(t)$, and set $Z_{i}(t)=Z_{i}(t-1)$ for all $i\neq I(t)$
end for
Set $A:=\{i:Z_{i}(m)\neq\boldsymbol{0}\}$
Define
$\widehat{P}_{i}:=\widehat{P}_{i}(m)=\begin{cases}\frac{Z_{i}(m)}{\Arrowvert Z_{i}(m)\Arrowvert_{1}}&\text{ for $i\in A$,}\\ P\in\Delta_{(d-1)}\text{ arbitrarily }&\text{ for $i\in A^{c}$}.\end{cases}$
$\widehat{\mu}_{n}(m):=\sum_{i=1}^{n}\frac{1}{n}\delta_{\widehat{P}_{i}}$.
Algorithm 2 Shallow sequencing$(m,\boldsymbol{P},\boldsymbol{u})$
In order to ensure the noisy empirical distribution is a good approximation, we need the assumption on $\boldsymbol{u}$. Recalling Algorithm 2, a cell is selected randomly according to $\boldsymbol{u}$ at each iteration. If most $u_{i}$’s are too small, for example $u_{i}=O(n^{-2})$, then it is unlikely to extract gene expression from those cells so that the noisy empirical distribution cannot be close to the truth. Therefore, it should be required that most of cells can be selected according to $\boldsymbol{u}$.
Assumption 2.3.
For any $n\in\mathbb{N}$, $\boldsymbol{u}$ is concentrated on $\{u\in\Delta_{(n-1)}:|\{i:u_{i}=o(n^{-1})|=o(n)\}$: more precisely, there exists a deterministic constant $c_{*}>0$ independent of $n$ such that for $n-o(n)$ many $u_{i}$’s, it holds almost surely that
$u_{i}\geq\frac{c_{*}}{n}.$ (2.2)
The first theorem answers to question (1): given an error tolerance $\varepsilon$ and $n$-cells, how many reads does Algorithm 2 require to guarantee that Wasserstein distance between the noisy empirical distribution and the true empirical one is less than $\varepsilon$.
Theorem 2.4.
Fix $p\in[1,2]$, and suppose that 2.3 holds. Then,
$\mathbb{E}W_{p}(\widehat{\mu}_{n}(m),\mu_{n})\leq\sqrt{\frac{8\mathbb{E}\left\|{P}\right\|_{0}}{c_{*}}\,\frac{n}{m}}+o_{n}(1).$ (2.3)
Additionally, given $\varepsilon>0$, $n\in\mathbb{N}$ and $\alpha\in(0,1)$, if $m$ satisfies
$m\geq\frac{8(1+\alpha)n\mathbb{E}\Arrowvert P\Arrowvert_{0}}{c_{*}\varepsilon^{2}}-\frac{8(1+\alpha)\varepsilon^{p-2}n\mathbb{E}\Arrowvert P\Arrowvert_{0}}{c_{*}}\sqrt{\frac{\alpha\log n}{n}},$ (2.4)
then it holds that
$W_{p}(\widehat{\mu}_{n}(m),\mu_{n})\leq\varepsilon$ (2.5)
with probability $1-O\left(n^{-\frac{\alpha}{2}}\right)$. Here the hidden constants in $O\left(n^{-\frac{\alpha}{2}}\right)$ depend on $\alpha$, $p$, $\mathbb{E}\Arrowvert P\Arrowvert_{0}$ and $\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}$.
Regarding (2.4), a lower bound of $m$ is proportional to $\alpha,\mathbb{E}\Arrowvert P\Arrowvert_{0},\varepsilon^{-1},c_{*}^{-1},p^{-1}$ (obviously, as $n$ larger, $m$ should be larger). $\alpha$ determines the rate of the probability of the event of interest at which the probability of the event $W_{p}(\widehat{\mu}_{n}(m),\mu_{n})>\varepsilon$ decays to 00 as $n\to\infty$. One can choose $\alpha$ such that it vanishes as $n$ goes $\infty$ while not too fast: for example $\alpha=O\left(\frac{\log\log n}{\log n}\right)$.
An interesting parameter is $\mathbb{E}\Arrowvert P\Arrowvert_{0}$, the average sparsity of $P$. Observing Algorithm 2, $E(t)$ is chosen according to $P_{I}$ at every iteration. If such $P_{I}$ is sparse, then a fewer iterations are sufficient to approximate $P_{I}$ well. As $n$ large, concentration phenomenon occurs so that most of $\Arrowvert P_{i}\Arrowvert_{0}$’s concentrate on $\mathbb{E}\Arrowvert P\Arrowvert_{0}$. For this reason, the average sparsity plays a role in order to determine a lower bound of the effective number of reads.
The inverse proportion of $\varepsilon$ is usual but the dependence of $c_{*}$ needs to be explained. Observing Algorithm 2 again, a cell $I$ is chosen according to $\boldsymbol{u}$ at every iteration. Gene expressions of $n-o(n)$ many cells should be measured enough to achieve (2.5). It is then necessarily satisfied that $n-o(n)$ many cells appear during running Algorithm 2. Unless $c_{*}$ is $O(1)$, the event that $O(n)$ many cells are not observed enough would happen with nonnegligible probability.
Theorem 2.4 tells us that $\mathbb{E}W_{p}(\widehat{\mu}_{n}(m),\mu_{n})\lesssim\sqrt{\frac{n}{m}}+o_{n}(1)$, suggesting that $m$ needs to grow faster than $n$ in order to obtain a diminishing Wasserstein distance between $\widehat{\mu}_{n}(m)$ and $\mu_{n}$. The following theorem provides a lower bound, confirming that this is indeed the case under certain assumptions.
Theorem 2.5.
Let $\mathrm{dist(\cdot,\cdot)}$ be the $\ell_{q}$ metric on $\Delta_{(d-1)}$ for $q\in[1,2]$, and assume uniform cell weights $\boldsymbol{u}=\left(\frac{1}{n},\dots,\frac{1}{n}\right)$. Then for any $p\geq 1$,
$\mathbb{E}W_{p}(\widehat{\mu}_{n}(m),\mu_{n})\geq\frac{1-\mathbb{E}\|P\|_{2}^{2}}{4}\frac{n}{m}$
provided $\mathbb{E}\|P\|_{2}^{2}<1$ and $m\geq 2n\log\frac{4}{1-\mathbb{E}\|P\|_{2}^{2}}$.
The next theorem and its corollary answer question (2), which is the main theme of this paper: given the number of $m$ reads, how many cells can one use in order to ensure that the noisy empirical distribution is as close as possible to the true distribution $\mu$.
Let us recall a trade-off for approximating $\mu$ in RNA sequencing: on the one hand a larger $n$ (i.e. more cells) helps get a better approximation of $\mu$ as usual because it allows to explore the population more, but on the other hand, since $m$ is fixed, if $n$ is too large compared to $m$, gene expression of each cell cannot be measured precisely well so that the noisy empirical distribution is poorly qualified. Therefore, the optimal number of cells should be determined sophisticatedly by taking into account these contradictory aspects.
Theorem 2.6.
Fix $\alpha\in(0,1)$ and $p\in[1,2]$. Assume that there are some $k>\max\{d^{*}_{p}(\mu),4\}$ and $m\geq m_{0}$ for some $m_{0}=m_{0}(c_{*},\alpha,\mathbb{E}\Arrowvert P\Arrowvert_{0})>0$.
Under Assumption 2.3, if $n$ satisfies (2.4), then Algorithm 2 $(m,\boldsymbol{P},\boldsymbol{u})$ outputs $\widehat{\mu}_{n}(m)$ which satisfies
$W_{p}(\widehat{\mu}_{n}(m),\mu)\leq O\left(n^{-\frac{1}{k}}\right)$ (2.6)
with probability at least $1-O\left(n^{-\frac{\alpha}{2}}\right)$. Here the hidden constants in $O\left(n^{-\frac{\alpha}{2}}\right)$ depend on $\alpha$, $p$, $\mathbb{E}\Arrowvert P\Arrowvert_{0}$ and $\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}$.
A parameter $k$ can be regarded as the intrinsic dimension of the support of $\mu$. In reality, the intrinsic dimension $k$ is much smaller than the ambient dimension. If the support of $\mu$ has a nice low dimensional structure, then given the same number of reads it is allowed to use more cells to approximate population more, which helps the noisy empirical distribution close to the truth. In words, the low dimension structure allows to leverage more cells to explore population more widely.
The intrinsic dimension is unknown a priori in reality. Possibly, it can be estimated by using a few of $\widehat{P}_{i}$’s provided by Algorithm 2 first, and then choose an optimal $n$ based on the estimation of the dimension. Notice that, however, any dimension estimator based on $\widehat{P}_{i}$’s underestimates the true one because $\widehat{P}_{i}$ is always sparser than $P_{i}$ due to the nature of sequencing. In our empirical analysis, we adopt PCA to estimate $k$: see Section 3.1.
One must be careful about the relation between the sparsity of $P$ and the intrinsic dimension of $\mu$. It is possible that even if $\mu$ has a small intrinsic dimension, $P$ is not sparse. Suppose $\mu$ is concentrated on the diagonal, the span of $\{(1,\dots,1)\}$. In this case $d^{*}_{p}(\mu)=1$ but $\Arrowvert P\Arrowvert_{0}=d$. On the other hand, however, if $\Arrowvert P\Arrowvert_{0}$ is concentrated at most $r\ll d$, then $\mu$ must exhibit a lower intrinsic dimensionality. In this sense, the sparsity of $P$ implies a lower intrinsic dimension of $\mu$, but the converse does not hold in general.
According to Theorem 2.6, one obtains the following corollary, which is straightforward from Theorem 2.6. For this reason, the proof is omitted. Here, we ignore lower order terms, and focus on the parameters of interest only.
Corollary 2.7.
Fix $\alpha\in(0,1)$ and $p\in[1,2]$. Assume that there are some $k>\max\{d^{*}_{p}(\mu),4\}$ and $m\geq m_{0}$ for some $m_{0}=m_{0}(c_{*},\alpha,\mathbb{E}\Arrowvert P\Arrowvert_{0})>0$. Under 2.3, if $n$ satisfies
$n\sim\left(\frac{m}{\mathbb{E}\Arrowvert P\Arrowvert_{0}}\right)^{1-\frac{2}{k+2}},$ (2.7)
then the output $\widehat{\mu}_{n}(m)$ by Algorithm 2 $(m,\boldsymbol{P},\boldsymbol{u})$ achieves
$W_{p}(\widehat{\mu}_{n}(m),\mu)\lesssim\left(\frac{\mathbb{E}\Arrowvert P\Arrowvert_{0}}{m}\right)^{\frac{1}{k+2}}$ (2.8)
with probability at least $1-O\left(\left(\frac{m}{\mathbb{E}\Arrowvert P\Arrowvert_{0}}\right)^{-\frac{\alpha}{3}}\right)$.
It is worthwhile to mention that the order (2.8) is almost optimal. If the support of $\mu$ has the intrinsic dimension $k$, then $W_{p}(\mu_{n},\mu)=O\left(n^{-\frac{1}{k}}\right)$ with high probability [Niles-Weed_Bach, Theorem 1, Proposition 20]. Since $n\sim m^{1-o_{k}(1)}$ and $W_{p}(\widehat{\mu}_{n}(m),\mu)\sim m^{-\frac{1}{k+2}}$, (2.8) achieves the almost optimal rate of convergence.
Notice that (2.7) almost matches the recent result in [zhang2020determining]. Our result is more general in the sense that we do not assume anything on $\mu$: in [zhang2020determining] the authors consider only exponential families.
We want to emphasize that (2.8) explicitly illustrates how the sparsity and the intrinsic dimension play a role in the optimal allocation, which the previous literature didn’t capture. It shows that Wasserstein distance increases as $k$ and $\mathbb{E}\Arrowvert P\Arrowvert_{0}$ increase, which perfectly coincides with our intuition. Trivially since larger $k$ and $\mathbb{E}\Arrowvert P\Arrowvert_{0}$ require to measure more genes, the Wasserstein distance cannot but increase whenever the number of reads is fixed.
[…]
Corollary 2.8.
Let $\mathrm{dist(\cdot,\cdot)}$ be the $\ell_{q}$ metric on $\Delta_{(d-1)}$ for $q\in[1,2]$, and assume uniform cell weights $\boldsymbol{u}=\left(\frac{1}{n},\dots,\frac{1}{n}\right)$. As before, let $k>\max\{d_{p}^{*}(\mu),4\}$. Then for any $p\geq 1$,
$\mathbb{E}W_{p}(\widehat{\mu}_{n}(m),\mu)\geq\left(\frac{1-\mathbb{E}\left\|{P}\right\|_{2}^{2}}{4}\frac{n}{m}-O\left(n^{-\frac{1}{k}}\right)\right)_{+}$ (2.9)
provided $\mathbb{E}\|P\|_{2}^{2}<1$ and $m\geq 2n\log\frac{4}{1-\mathbb{E}\|P\|_{2}^{2}}$. We have used the notation $t_{+}:=\max(t,0)$.
Remark 2.9.
In order to obtain a non-trivial lower bound of (2.9),
$n\sim m^{1-\frac{1}{k}}.$
Recalling the exponent of (2.7), an upper bound (2.8) is achieved by
$n\sim m^{1-\frac{2}{k+2}}.$
For $k\geq 2$, $1-\frac{2}{k+2}$ is always larger than $1-\frac{1}{k}$. In this sense, although both exponents are asymptotically the same as $k\to\infty$, there is a small gap in the exponents of upper bound and lower bound of $n$ in terms of $m$ for finite $k$. It would be interesting to match this gap but we leave it for future work.
3. Empirical results
3.1. Experimental setup
In this section, we empirically investigate the dependence of the optimal number of samples $n^{*}$ on the sequencing budget $m$ via simulation experiments on a real single-cell dataset. Given a ground truth population distribution $\mu$ constructed from a real dataset, we generate $n$ i.i.d samples followed by $m$ reads for various values of $n$ and $m$, yielding noisy empirical distributions $\hat{\mu}_{n}(m)$ (recall the sampling and sequencing procedures from Section 2.1). The resulting Wasserstein errors $W_{1}(\hat{\mu}_{n}(m),\mu)$ then reveal the desired dependence, demonstrating the validity of our theoretical results from Section 2.2.
Dataset
[…]
$\mathbf{C}=\begin{bmatrix}-&C_{1}^{T}&-\\ &\vdots&\\ -&C_{N}^{T}&-\end{bmatrix}\,\in\mathbb{N}_{0}^{N\times d},$
whose rows correspond to the $N$ cells measured. Note that although this matrix was obtained experimentally, which necessarily corresponds to an ‘empirical’ distribution in reality, we treat it as the ground truth population in our simulations. Particularly, we construct the distribution $\mu$ as
$\mu=\frac{1}{N}\sum_{\ell=1}^{N}\delta_{C_{\ell}/\|C_{\ell}\|_{1}}\in\mathcal{P}(\Delta_{(d-1)}).$ (3.1)
Note that $\mu$ defined in such a manner is an atomic measure; its discrete nature enables easy generation of (pseudorandom) samples and computation of Wasserstein distances $W_{1}(\hat{\mu}_{n}(m),\mu)$. For $\mu$ constructed from the reprogramming dataset considered in this section, we have $d=1000$, $N=4851$ and $\mathbb{E}_{P\sim\mu}\|P\|_{0}\simeq 175$.
Cell weights
Recall that the noisy sequencing process (Algorithm 2) depends on the normalized cell weights $\boldsymbol{u}=(u_{1},\dots,u_{n})\in\Delta_{n-1}$, which, as discussed earlier, may or may not be independent of the gene expression vectors $\{P_{1},\dots,P_{n}\}$. A simple scenario is that of uniform cell weights $\boldsymbol{u}=\left(\frac{1}{n},\dots,\frac{1}{n}\right)$, which we investigate first (Section 3.2, Section 3.3). A general formulation considers the joint distribution $(P,U)\sim\mu_{(P,U)}$, and this is studied in Section 3.4.
Assessing the intrinsic dimension
The theoretical results from Section 2.2 depend on the intrinsic dimension (specifically, the upper Wasserstein dimension) of $\mu$. A natural heuristic for approximating this quantity is the dimension of an affine subspace in $\mathbb{R}^{d}$ that captures most of the variance of $\mu$. This can be computed conveniently via principal component analysis (PCA); we set the approximate intrinsic dimension $k$ to be the smallest number of principal components that cumulatively capture at least 95% of the variance of $\mu$. For the reprogramming dataset, this procedure gives $k=83$ (which is significantly smaller than the ambient dimension $d=1000$); see Figure A.1.
A key part of our results from Section 2.2 is the influence of $k$ on the $m$-$n$ relationship. We validate this further in Section 3.3 by conducting tests on lower-dimensional versions $\mu$ that we construct via non-negative matrix factorization of the normalized counts matrix.
3.2. Results on the reprogramming dataset with uniform cell weights
For various pairs of $n$ and $m$, we simulate the noisy empirical distribution $\widehat{\mu}_{n}(m)$ for ten (10) independent trials, and compute the mean Wasserstein error $W_{1}(\widehat{\mu}_{n}(m),\mu)$ over those trials. Contours of mean $W_{1}(\widehat{\mu}_{n}(m),\mu)$ over $m$-$n$ axes are plotted in Figure 2. For each budget $m$ tested, we identify the number of cells $n=n^{*}$ that yields the lowest mean $W_{1}(\widehat{\mu}_{n}(m),\mu)$. This $m$-$n^{*}$ relationship, shown in Figure 2, appears to be well captured by our theoretically derived optimal allocation
$n\sim\left(\frac{Cm}{\mathbb{E}\|P\|_{0}}\right)^{1-\frac{2}{2+k}}$
where $C$ is a $\Theta(1)$ constant. In the plots, we have used the value $C=1/2$ for this dataset. The exponent $1-\frac{2}{2+k}$ shows up as the slope on the log-scale plot Figure 2 (right panel).
Figure 2. Optimal allocation ($m$-$n$ relationship) with uniform cell weights. The contours and the colour scale show $W_{1}(\widehat{\mu}_{n}(m),\mu)$ averaged over ten (10) independent realizations of $\widehat{\mu}_{n}(m)$, for various pairs of $m$ and $n$ on a grid. The $m$ and $n$ axes are plotted in a linear scale on the left panel, and in a log scale on the right panel. The optimal values of $n^{*}$ for each $m$, as observed in this simulation, are plotted with blue solid dots. Our theoretically derived optimal allocation is plotted as an orange dash-dot line.
3.3. Results for various intrinsic dimensions
The optimal allocation is crucially dependent on the intrinsic dimension $k$. We investigate this further by constructing synthetic versions of the reprogramming dataset with different values of $k$, to produce synthetic population distributions $\mu_{k}$. Details on these constructions, and some statistics of the synthetic datasets, are presented in Section A.3. We repeat the experiment from Section 3.2 on the synthetic distributions $\mu_{k}$ for various values of $k$; the results are presented in Figure 3. As before, the observed $\log m$ $\log n^{*}$ relationships are reasonably approximated by the respective theoretically derived optimal allocations (with $C=1/2$).
Figure 3. Optimal allocations for different intrinsic dimensions $k$. Each panel shows the optimal allocation results for a synthetic population distribution $\mu_{k}$ of a different intrinsic dimension $k$. The contours and the colour scale show $W_{1}(\widehat{\mu}_{n}(m),\mu_{k})$ averaged over ten (10) independent realizations of $\widehat{\mu}_{n}(m)$, for various pairs of $m$ and $n$. The optimal $\log n^{*}$ for each $\log m$, as observed in the simulations, are plotted with blue solid dots. Our theoretically derived optimal allocations are plotted as orange dash-dot lines.
3.4. Results in the case of non-uniform cell weights
To simulate sequencing with general, non-uniform cell weights, we construct the ground truth distribution $\mu_{(P,U)}$ from real data as before. Since the number of reads that a cell receives is an indication of its sampling frequency, we can use the row sums of $\mathbf{C}$ to help construct $\mu_{(P,U)}$. In order to capture the possible dependence or independence between $P$ and $U$, two scenarios are considered in our experiments.
Scenario 1
$U$ dependent on $P$. Here, we always match a sampled cell to its original weight in the ground truth counts matrix $\mathbf{C}$. More precisely,
$\mu_{(P,U)}=\frac{1}{N}\sum_{\ell=1}^{N}\delta_{(C_{\ell}/\|C_{\ell}\|_{1}\,,\,\|C_{\ell}\|_{1})}.$
Scenario 2
$U$ independent of $P$. Here, a sampled cell is paired with an independently chosen weight from the ground truth counts matrix $\mathbf{C}$:
$\mu_{(P,U)}=\frac{1}{N^{2}}\sum_{\ell,k=1}^{N}\delta_{(C_{\ell}/\|C_{\ell}\|_{1}\,,\|C_{k}\|_{1})}.$
Note that the two scenarios only differ in their couplings of $P$ and $U$; the marginals of $P$ and $U$ remain the same. The $P$-marginal in both cases equals $\mu$ as defined in (3.1).
Our optimal allocation theory is agnostic to the possible dependence between $P$ and $U$, and we verify that this is also the case in simulations. We repeat the experiment from Section 3.2 under scenarios 1 and 2, and present the results in Figure 4 and Figure 5 respectively. The behavior is similar to that seen in Figure 2; the observed $m$-$n^{*}$ relationships are reasonably close to the theoretically derived optimal allocations with $C=1/2$. Also, considering the effect of the parameter $c_{*}$ in Theorem 2.4, readjusting $C$ can result in an even better fit between the theory and simulations.
Figure 4. Optimal allocation with non-uniform cell weights: Scenario 1 ($U$ dependent on $P$). The contours and the colour scale show $W_{1}(\widehat{\mu}_{n}(m),\mu)$ averaged over ten (10) independent realizations of $\widehat{\mu}_{n}(m)$, for various pairs of $m$ and $n$. The $m$ and $n$ axes are plotted in a linear scale on the left panel, and in a log scale on the right panel. The optimal values of $n^{*}$ for each $m$, as observed in this simulation, are plotted with blue solid dots. Our theoretically derived optimal allocation is plotted as an orange dash-dot line.
Figure 5. Optimal allocation with non-uniform cell weights: Scenario 2 ($U$ independent of $P$). The contours and the colour scale show $W_{1}(\widehat{\mu}_{n}(m),\mu)$ averaged over ten (10) independent realizations of $\widehat{\mu}_{n}(m)$, for various pairs of $m$ and $n$. The $m$ and $n$ axes are plotted in a linear scale on the left panel, and in a log scale on the right panel. The optimal values of $n^{*}$ for each $m$, as observed in this simulation, are plotted with blue solid dots. Our theoretically derived optimal allocation is plotted as an orange dash-dot line.
4. Proofs
Let’s recall the problem setting. $\boldsymbol{P}:=\{P_{1},\dots,P_{n}\}\subseteq\Delta_{(d-1)}$ and $\boldsymbol{u}=(u_{1},\dots,u_{n})\in\Delta_{(n-1)}$ denote the set of gene expressions of $n$-cells and the vector of the weight vector of cells, respectively, and $\mu$ is the unknown ground truth distribution. A true empirical distribution and the noisy one obtained by Algorithm 2 are denoted by
$\mu_{n}:=\sum_{i=1}^{n}\frac{1}{n}\delta_{P_{i}}\text{ and }\widehat{\mu}_{n}(m):=\sum_{i=1}^{n}\frac{1}{n}\delta_{\widehat{P}_{i}},$
respectively.
4.1. Proof of Theorem 2.4
The first objective is to control $W_{p}(\widehat{\mu}_{n}(m),\mu_{n})$. The following lemma allows to bound $W_{p}(\widehat{\mu}_{n}(m),\mu_{n})$ by $\left(\sum\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\right)^{\frac{1}{p}}$.
Lemma 4.1.
[Oldandnew, Theorem 4.8] Let $\mathcal{X}$ and $\mathcal{Y}$ be Polish spaces, and $c:\mathcal{X}\times\mathcal{Y}\longrightarrow\mathbb{R}\cup\{\infty\}$ be a lower semicontinuous function. For each $\mu\in\mathcal{P}(\mathcal{X})$ and $\nu\in\mathcal{P}(\mathcal{Y})$
[…]
the associated optimal transport cost functional. Let $(\Theta,\lambda)$ be a probability space, and let $\mu_{\theta},\nu_{\theta}$ be two measurable probability measures on $\mathcal{P}(\mathcal{X})$ and $\mathcal{P}(\mathcal{Y})$, respectively. Assume that $c(x,y)\geq a(x)+b(y)$, where $a\in L^{1}(d\mu_{\theta}d\lambda(\theta))$ and $b\in L^{1}(d\nu_{\theta}d\lambda(\theta))$. Then,
[…]
Lemma 4.1 states the convexity of optimal transport functional. Since norms over $\mathbb{R}^{d}$ are lower semicontinuous, Jensen’s inequality implies that $W_{p}$ for any $p\geq 1$ enjoys convexity. More precisely,
$W^{p}_{p}(\widehat{\mu}_{n}(m),\mu_{n})\leq\sum_{i=1}^{n}\frac{1}{n}W^{p}_{p}(\delta_{\widehat{P}_{i}},\delta_{P_{i}})=\sum_{i=1}^{n}\frac{1}{n}\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}.$
Hence, one gets
$W_{p}(\widehat{\mu}_{n}(m),\mu_{n})\leq\left(\sum_{i=1}^{n}\frac{1}{n}\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\right)^{\frac{1}{p}}.$ (4.1)
Therefore, the problem reduces to achieving $\sum\frac{1}{n}\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}$ to be small with high probability.
$\sum\frac{1}{n}\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}$ would be small by concentration phenomenon if $n$ is large, and also $m$ is large enough proportional to $n$. The concentration inequality of the estimator of multinomial distribution is well-known. The next lemma is regarding the concentration inequality of multinomial distribution. Originally, [weissman2003inequalities, Theorem 2.1] concerns only $\ell_{1}$ norm. However, since $\Arrowvert P-P^{\prime}\Arrowvert_{1}\geq\Arrowvert P-P^{\prime}\Arrowvert_{q}$ holds on $\Delta_{(d-1)}$ for any $1\leq q\leq\infty$, one can extend the concentration inequality to $\ell_{q}$ norms. Recalling the $\ell_{0}$ norm defined in (2.1), the concentration inequality depends on an error parameter $\delta$ and $\Arrowvert P\Arrowvert_{0}$, the sparsity of $P$, and the number of iteration $T$.
Lemma 4.2.
[weissman2003inequalities, Theorem 2.1][qian2020concentration, Proposition 1] Let $P\in\Delta_{(d-1)}$ and $\widetilde{P}=\widetilde{P}(T)$ be the output of Algorithm 1$(T,P)$. Then, for any $\delta\in[0,1]$ and $p\in[1,\infty)$,
$\mathbb{P}\left(\mathrm{dist}(\widetilde{P},P)^{p}\leq\left(\frac{2\Arrowvert P\Arrowvert_{0}\log\left(\frac{1}{\delta}\right)}{T}\right)^{\frac{p}{2}}\right)\geq 1-\delta.$ (4.2)
Let $\varepsilon\in(0,1)$ be the fixed error tolerance and $\boldsymbol{T}:=(T_{1},\dots,T_{n})$ be the vector of the numbers of appearance of cells in Algorithm 2, i.e., for $1\leq i\leq n$
$T_{i}:=\sum_{t=1}^{m}\mathds{1}_{I(t)=i}.$ (4.3)
$\boldsymbol{T}$ is a multinomial vector with $m$ trials, $n$ integers and a probability vector $\boldsymbol{u}$. Notice that each marginal $T_{i}$ follows binomial distribution with $m$ trials and a success probability $u_{i}$.
Let $\alpha_{i}:=\exp\left(-\frac{T_{i}\varepsilon^{2}}{2\Arrowvert P_{i}\Arrowvert_{0}}\right)$. (4.2) leads to for each $i$ and $k\in\{1,\dots,\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor\}$
$\mathbb{P}\left(\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\geq k\varepsilon^{p}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)\leq\alpha_{i}^{k^{\frac{2}{p}}}.$
In particular,
$\mathbb{P}\left((k-1)\varepsilon^{p}\leq\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}<k\varepsilon^{p}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)\leq\alpha_{i}^{(k-1)^{\frac{2}{p}}}-\alpha_{i}^{k^{\frac{2}{p}}}.$
Here $\alpha_{i}$ is a random variable depending on $T_{i}$ and $\boldsymbol{P},\boldsymbol{u}$. Notice that $\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\leq 2^{p}$ almost surely.
A next lemma is regarding the construction of a new random variable $X_{i}$ which stochastically dominates $\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}$.
Lemma 4.3.
For each $1\leq i\leq n$ define a random variable $X_{i}$ which takes a value $\{\varepsilon^{p},\dots,\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor\varepsilon^{p},2^{p}\}$ as
$\mathbb{P}\left(X_{i}=k\varepsilon^{p}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)=\begin{cases}\alpha_{i}^{(k-1)^{\frac{2}{p}}}-\alpha_{i}^{k^{\frac{2}{p}}}&\text{if }k=1,\dots,\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor,\\ \alpha_{i}^{\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor^{\frac{2}{p}}}&\text{if }k=\frac{2^{p}}{\varepsilon^{p}}.\end{cases}$ (4.4)
Then $X_{i}$ stochastically dominates $\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}$ conditioned on $\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}$.
Proof.
One needs to show that for any $t\in[0,2^{p}]$
$\mathbb{P}\left(X_{i}\geq t\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)\geq\mathbb{P}\left(\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\geq t\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)$
and the strict inequality holds for some $t$ unless the distribution of $\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}$ is equal to that of $X_{i}$ conditioned on $\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}$. From (4.2) and (4.4), for $t\in((k-1)\varepsilon^{p},k\varepsilon^{p})$ it is straightforward that
$\mathbb{P}\left(\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\geq t\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)$ $\leq\mathbb{P}\left(\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\geq(k-1)\varepsilon^{p}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)$
$\leq\alpha_{i}^{(k-1)^{\frac{2}{p}}}$
$=\mathbb{P}\left(X_{i}\geq(k-1)\varepsilon^{p}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right).$
It is immediate that $\sum X_{i}$ stochastically dominates $\sum\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}$ by the following lemma.
Lemma 4.4.
[roch2024modern, Corollary 4.2.7] If $X_{i}$ stochastically dominates $Y_{i}$ for $i=1,2$, then $X_{1}+X_{2}$ stochastically dominates $Y_{1}+Y_{2}$.
One advantage of the construction of $X_{i}$’s is that it is easy to compute an upper bound of the expectation of $X_{i}$.
Lemma 4.5.
Let $X_{i}$ is a random variable defined in Lemma 4.3. Let $p\in[1,2]$. Then,
$\mathbb{E}\left[X_{i}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right]\leq\left(\frac{\varepsilon^{p-2}8n\Arrowvert P_{i}\Arrowvert_{0}}{c_{*}m}\right)\wedge 2^{p}$ (4.5)
where $x\wedge y:=\min\{x,y\}$.
Proof.
A straightforward calculation yields
$\mathbb{E}\left[X_{i}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right]$ $=\sum_{k=1}^{\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor}k\varepsilon^{p}\left(\alpha_{i}^{(k-1)^{\frac{2}{p}}}-\alpha_{i}^{k^{\frac{2}{p}}}\right)+2^{p}\alpha_{i}^{\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor^{\frac{2}{p}}}$
$=\varepsilon^{p}-\varepsilon^{p}\alpha_{i}+2\varepsilon^{p}\alpha_{i}-2\varepsilon^{p}\alpha_{i}^{2^{\frac{2}{p}}}+3\varepsilon^{p}\alpha_{i}^{2^{\frac{2}{p}}}-3\varepsilon^{p}\alpha_{i}^{3^{\frac{2}{p}}}+\dots+2^{p}\alpha_{i}^{\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor^{\frac{2}{p}}}$
$\leq\varepsilon^{p}\sum_{k=0}^{\lfloor\frac{2^{p}}{\varepsilon^{p}}\rfloor}\alpha_{i}^{k^{\frac{2}{p}}}$
$\leq\varepsilon^{p}\sum_{k=0}^{\infty}\alpha_{i}^{k^{\frac{2}{p}}}$
$\leq\varepsilon^{p}\sum_{k=0}^{\infty}\alpha_{i}^{k}$
where the last inequality follows from the fact that $\alpha_{i}<1$ and $1\leq p\leq 2$. Of course, the above infinite sum is useful only if it is less than $2^{p}$, which is the trivial upper bound. Hence, we have the following upper bound of the conditional expectation of $X_{i}$:
$\mathbb{E}\left[X_{i}\,|\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right]\leq\left(\varepsilon^{p}\sum_{k=0}^{\infty}\alpha_{i}^{k}\right)\wedge 2^{p}.$
Let $\gamma_{i}:=\exp\left(-\frac{\varepsilon^{2}}{2\Arrowvert P_{i}\Arrowvert_{0}}\right)$ for simplicity. Since each $T_{i}$ is the binomial random variable with parameters $m$ and $u_{i}$, it follows from the moment generating function of $T_{i}$ that
$\mathbb{E}\left[\alpha_{i}\,|\,\boldsymbol{P},\boldsymbol{u}\right]=\mathbb{E}\left[\gamma_{i}^{T_{i}}\,|\,\boldsymbol{P},\boldsymbol{u}\right]=\left((1-u_{i})+u_{i}\gamma_{i}\right)^{m}=\left(\left(1-\left(1-\gamma_{i}\right)u_{i}\right)^{\frac{1}{u_{i}}}\right)^{u_{i}m}.$
Since $(1-\frac{1}{x})^{x}$ monotonically increases and converges to $e^{-1}$, provided that $u_{i}m=O(1)$, one obtains
$\mathbb{E}\left[\alpha_{i}\,|\,\boldsymbol{P},\boldsymbol{u}\right]\leq\exp\left(-u_{i}m\left(1-\gamma_{i}\right)\right).$
One gets similarly for $k\geq 1$
$\mathbb{E}\left[\alpha_{i}^{k}\,|\,\boldsymbol{P},\boldsymbol{u}\right]=\mathbb{E}\left[(\gamma_{i}^{k})^{T_{i}}\,|\,\boldsymbol{P},\boldsymbol{u}\right]\leq\exp\left(-u_{i}m\left(1-\gamma_{i}^{k}\right)\right).$
Since for $k\geq 1$
$1-\gamma_{i}^{k}=1-\exp\left(-\frac{k\varepsilon^{2}}{2\Arrowvert P_{i}\Arrowvert_{0}}\right)=\frac{k\varepsilon^{2}}{2\Arrowvert P_{i}\Arrowvert_{0}}-O\left(\left(\frac{k\varepsilon^{2}}{2\Arrowvert P_{i}\Arrowvert_{0}}\right)^{2}\right)\geq\frac{k\varepsilon^{2}}{4\Arrowvert P_{i}\Arrowvert_{0}},$
an upper bound of $\mathbb{E}\left[\alpha_{i}^{k}\,|\,\boldsymbol{P},\boldsymbol{u}\right]$ is then given as
$\mathbb{E}\left[\alpha_{i}^{k}\,|\,\boldsymbol{P},\boldsymbol{u}\right]\leq\exp\left(-u_{i}m\left(1-\gamma_{i}^{k}\right)\right)\leq\exp\left(-u_{i}m\frac{k\varepsilon^{2}}{4\Arrowvert P_{i}\Arrowvert_{0}}\right).$
Thus, an upper bound of the conditional expectation of $X_{i}$ is derived as
$\mathbb{E}\left[X_{i}\,|\,\boldsymbol{P},\boldsymbol{u}\right]$ $\leq\left(\varepsilon^{p}\sum_{k=0}^{\infty}\mathbb{E}\left[\alpha_{i}^{k}\,|\,\boldsymbol{P},\boldsymbol{u}\right]\right)\wedge 2^{p}$
$\leq\left(\varepsilon^{p}\sum_{k=0}^{\infty}\exp\left(-u_{i}m\frac{k\varepsilon^{2}}{4\Arrowvert P_{i}\Arrowvert_{0}}\right)\right)\wedge 2^{p}$
$=\left(\frac{\varepsilon^{p}}{1-\exp\left(-u_{i}m\frac{\varepsilon^{2}}{4\Arrowvert P_{i}\Arrowvert_{0}}\right)}\right)\wedge 2^{p}$
$\leq\left(\frac{\varepsilon^{p-2}8n\Arrowvert P_{i}\Arrowvert_{0}}{c_{*}m}\right)\wedge 2^{p}.$
Combining Lemmas 4.3 and 4.4, we get
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathrm{dist}(\widehat{P}_{i},P_{i})^{p}\geq t\,\bigg{|}\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)\leq\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq t\,\bigg{|}\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right).$ (4.6)
Although the expectations of $\mathrm{dist}(\widehat{P}_{i},P_{i})$’s are complicated, we now know an upper bound of the expectations of $X_{i}$’s. Hence, the strategy that we adopt is to derive the concentration inequality of the sum of $X_{i}$’s by using (4.5). Then combined with (4.6), the concentration will give a lower bound of $m$ given $n$ to achieve the error rate at most $\varepsilon$.
From (4.5), an upper bound of the conditional expectation of the sum of $X_{i}$’s is derived as
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\,\bigg{|}\,\boldsymbol{P},\boldsymbol{u}\right]\leq\frac{8\varepsilon^{p-2}}{c_{*}m}\sum_{u_{i}:u_{i}\geq\frac{c_{*}}{n}}\Arrowvert P_{i}\Arrowvert_{0}+o_{n}(1)$ (4.7)
provided that $m=O(n)$. It remains to bound the right hand side of (4.7) by $\varepsilon^{p}+o_{n}(1)$ with satisfying $m=O(n)$. A sufficient condition for it is
$m\geq\frac{8\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}}{c_{*}\varepsilon^{2}},$
which is $O(n)$ but still a random variable depending on $\Arrowvert P_{i}\Arrowvert_{0}$’s. To obtain a deterministic lower bound of $m$, we leverage the fact that $\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}$ also exhibits a concentration phenomenon.
Lemma 4.6.
Fix $\alpha\in(0,1)$. Then,
$\mathbb{P}\left(\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}\geq(1+\alpha)n\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)\leq\exp\left(-\frac{2\alpha^{2}(\mathbb{E}\Arrowvert P\Arrowvert_{0})^{2}n}{\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}}\right).$ (4.8)
Proof.
Notice that each $\Arrowvert P_{i}\Arrowvert_{0}$ is bound from 00 to $d$. Let $c_{j}$ be the maximum value of the $j$-th entry of $P$. Since $c_{j}\leq 1$, (2.1) leads to $\mu$-almost surely
$\sum_{j=1}^{d}c_{j}^{2}\leq\sup\left\{\sum_{j=1}^{d}\mathds{1}_{p_{j}\neq 0}:\mu\big{(}\{P\in\Delta_{(d-1)}:\sum_{j=1}^{d}\mathds{1}_{p_{j}\neq 0}>0\}\big{)}>0\right\}=\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}.$
Using Hoeffding’s inequality or McDiarmid’s inequality straightforwardly with the above bound implies that
$\mathbb{P}\left(\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}-\mathbb{E}\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}\geq t\right)\leq\exp\left(-\frac{2t^{2}}{n\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}}\right).$
Fix $\alpha\in(0,1)$. Choosing $t=\alpha\mathbb{E}\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}=\alpha n\mathbb{E}\Arrowvert P\Arrowvert_{0}$, it follows that
$\mathbb{P}\left(\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}\geq(1+\alpha)n\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)\leq\exp\left(-\frac{2\alpha^{2}(\mathbb{E}\Arrowvert P\Arrowvert_{0})^{2}n}{\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}}\right).$
Remark 4.7.
Since $\frac{(\mathbb{E}\Arrowvert P\Arrowvert_{0})^{2}}{\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}}$ does not depend on $n$, the right hand side of (4.8) decays as $n\to\infty$.
Now, we are ready to prove Theorem 2.4.
Proof of Theorem 2.4.
A direct consequence of (4.7) is that under 2.3,
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\right]\leq\frac{8n\mathbb{E}\left\|{P}\right\|_{0}}{\varepsilon^{2-p}c_{*}m}+o_{n}(1).$
Setting $\varepsilon^{2}=\frac{8n\mathbb{E}\left\|{P}\right\|_{0}}{c_{*}m}$, and noting that
$\mathbb{E}W_{p}(\widehat{\mu}_{n}(m),\mu_{n})\leq\left(\mathbb{E}W_{p}^{p}(\widehat{\mu}_{n}(m),\mu_{n})\right)^{\frac{1}{p}}\leq\left(\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\right]\right)^{\frac{1}{p}}$
concludes the bound in expectation (2.3).
If $m$ satisfies (2.4), combinied (4.7) and (4.8), it follows that
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\,\bigg{|}\,\boldsymbol{P},\boldsymbol{u}\right]\leq\varepsilon^{p}-\sqrt{\frac{\alpha\log n}{n}}+o_{n}(1).$
Notice that this upper bound is independent of $\boldsymbol{P},\boldsymbol{u}$. Therefore, if $m$ satisfies (2.4), then
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\right]=\mathbb{E}\left[\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\,\bigg{|}\,\boldsymbol{P},\boldsymbol{u}\right]\right]\leq\varepsilon^{p}-\sqrt{\frac{\alpha\log n}{n}}+o_{n}(1).$ (4.9)
To obtain an upper bound of the tail probability of $\frac{1}{n}\sum X_{i}$, since $X_{i}\in[0,2^{p}]$ almost surely, applying Hoeffding’s inequality with (4.9) leads to
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq\varepsilon^{p}\right)$
$=\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\right]\geq\varepsilon^{p}-\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\right]\right)$
$\leq\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}-\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}X_{i}\right]\geq\sqrt{\frac{\alpha\log n}{n}}-o_{n}(1)\,\Bigg{|}\,\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}\leq(1+\alpha)n\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)$
×$\quad\times\mathbb{P}\left(\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}\leq(1+\alpha)n\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)+\mathbb{P}\left(\sum_{i=1}^{n}\Arrowvert P_{i}\Arrowvert_{0}\geq(1+\alpha)n\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)$
$\leq\exp\left(-\frac{n\left(\sqrt{\frac{\alpha\log n}{n}}-o_{n}(1)\right)^{2}}{2^{2p-1}}\right)\left(1-\exp\left(-\frac{2\alpha^{2}(\mathbb{E}\Arrowvert P\Arrowvert_{0})^{2}n}{\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}}\right)\right)$
$\quad+\exp\left(-\frac{2\alpha^{2}(\mathbb{E}\Arrowvert P\Arrowvert_{0})^{2}n}{\operatorname*{ess\,sup}\Arrowvert P\Arrowvert_{0}}\right)$
$\leq O\left(n^{-\frac{\alpha}{2}}\right).$
Combining this with the stochastic domination relation (4.6), it follows that if $m$ satisfies (2.4),
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathrm{dist}(\widehat{P}_{i},P_{i})\geq\varepsilon^{p}\right)$ $=\mathbb{E}\left[\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathrm{dist}(\widehat{P}_{i},P_{i})\geq\varepsilon^{p}\,\bigg{|}\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)\right]$
$\leq\mathbb{E}\left[\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq\varepsilon^{p}\,\bigg{|}\,\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right)\right]$
$=\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq\varepsilon^{p}\right)$
$\leq O\left(n^{-\frac{\alpha}{2}}\right).$
Therefore, the conclusion follows from (4.1). ∎
4.2. Proof of Theorem 2.5
Since $W_{p}\geq W_{p^{\prime}}$ for $p\geq p^{\prime}$ by Jensen’s inequality [Oldandnew, Remark 6.6], it suffices to lower bound $W_{1}$. By Kantorovich duality [villani2003topics, Theorem 1.14], we have that
[…]
for any function $f:\Delta_{(d-1)}\to\mathbb{R}$ that is 1-Lipschitz with respect to $\mathrm{dist}(\cdot,\cdot)$ on $\Delta_{(d-1)}$. We choose $f(P)=\frac{\|P\|_{2}^{2}}{2}$, and verify its Lipschitz property: for any $P,P^{\prime}\in\Delta_{(d-1)}$ and $q\in[1,2]$,
$|f(P^{\prime})-f(P)|$ $=\frac{1}{2}\left|\|P^{\prime}\|_{2}^{2}-\|P\|_{2}^{2}\right|$
$=\frac{1}{2}\left(\|P^{\prime}\|_{2}+\|P\|_{2}\right)\left|\|P^{\prime}\|_{2}-\|P\|_{2}\right|$
$\leq\frac{1}{2}\cdot 2\cdot\|P^{\prime}-P\|_{2}\leq\|P^{\prime}-P\|_{q}.$
Now, recall that $\mu_{n}=\frac{1}{n}\sum\delta_{P_{i}}$ and $\widehat{\mu}_{n}(m)=\frac{1}{n}\sum\delta_{\widehat{P}_{i}}$ where $P_{1},\dots,P_{n}\overset{\mathrm{iid}}{\sim}\mu$ and $\widehat{P}_{1},\dots,\widehat{P}_{n}$ are generated as per Algorithm 2. Furthermore, recall $\boldsymbol{T}=(T_{1},\dots,T_{n})$ from (4.3) which is multinomially distributed as
$\boldsymbol{T}\,|\,\boldsymbol{u}\sim\mathrm{Multinomial}(m,\boldsymbol{u}),$
where we have assumed $\boldsymbol{u}=\left(\frac{1}{n},\dots,\frac{1}{n}\right)$. The fixed $\boldsymbol{u}$ also implies that $\boldsymbol{T}$ and $\boldsymbol{P}$ are independent. Putting these definitions together, we can write
$\mathbb{E}W_{1}(\widehat{\mu}_{n}(m),\mu_{n})$
[…]
$=\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\left(\frac{1}{2}\|\widehat{P}_{i}\|_{2}^{2}-\frac{1}{2}\|P_{i}\|_{2}^{2}\right)$
$=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{2}\underbrace{\mathbb{E}\mathds{1}_{\{T_{i}>0\}}\left(\|\widehat{P}_{i}\|_{2}^{2}-\|P_{i}\|_{2}^{2}\right)}_{\mathrm{(I)}}+\frac{1}{n}\sum_{i=1}^{n}\frac{1}{2}\underbrace{\mathbb{E}\mathds{1}_{\{T_{i}=0\}}\left(\|\widehat{P}_{i}\|_{2}^{2}-\|P_{i}\|_{2}^{2}\right)}_{\mathrm{(II)}},$
so that the problem simplifies to lower bounding the terms labeled (I) and (II).
We start with the simpler term (II). As defined in Algorithm 2, $\widehat{P}_{i}$ may be chosen arbitrarily in $\Delta_{(d-1)}$ when $T_{i}=0$, and hence, we use the trivial lower bound $\|\widehat{P}_{i}\|_{2}^{2}-\|P_{i}\|_{2}^{2}\geq-1$. Using $T_{i}\sim\mathrm{Binomial}(m,\frac{1}{n})$, this can be followed up with the estimate
$\mathbb{E}\mathds{1}_{\{T_{i}=0\}}=\mathbb{P}(T_{i}=0)=\left(1-\frac{1}{n}\right)^{m}\leq\exp\left(-\frac{m}{n}\right)$
to conclude that $(\mathrm{II})\geq-\exp\left(-\frac{m}{n}\right)$.
Next, we consider the term (I), rewriting it as
$\mathbb{E}\mathds{1}_{\{T_{i}>0\}}\left(\|\widehat{P}_{i}\|_{2}^{2}-\|P_{i}\|_{2}^{2}\right)=\mathbb{E}\mathds{1}_{\{T_{i}>0\}}\,\mathbb{E}\left[\|\widehat{P}_{i}\|_{2}^{2}-\|P_{i}\|_{2}^{2}\bigm{|}\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right].$
Using the fact that
$\widehat{P}_{i}\mid\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\sim\frac{1}{T_{i}}\mathrm{Multinomial}(T_{i},P_{i})$
whenever $T_{i}>0$, one can write the moments
$\mathbb{E}\left[\widehat{P}_{i}\bigm{|}\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right]$ $=P_{i},$
$\mathrm{tr}\,\left(\mathrm{Cov}\left[\widehat{P}_{i}\bigm{|}\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right]\right)$ $=\frac{1-\|P_{i}\|_{2}^{2}}{T_{i}},$
so that
$\mathbb{E}\left[\|\widehat{P}_{i}\|_{2}^{2}\bigm{|}\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right]=\|P_{i}\|_{2}^{2}+\frac{1}{T_{i}}(1-\|P_{i}\|_{2}^{2}).$
Hence, we simplify
$\mathbb{E}\left[\mathds{1}_{\{T_{i}>0\}}\,\mathbb{E}\left[\|\widehat{P}_{i}\|_{2}^{2}-\|P_{i}\|_{2}^{2}\bigm{|}\boldsymbol{T},\boldsymbol{P},\boldsymbol{u}\right]\right]$ $=\mathbb{E}\left[\mathds{1}_{\{T_{i}>0\}}\frac{1}{T_{i}}(1-\|P_{i}\|_{2}^{2})\right]$
$=\mathbb{E}\left[\mathds{1}_{\{T_{i}>0\}}\frac{1}{T_{i}}\right](1-\mathbb{E}\|P_{i}\|_{2}^{2})$
using the independence of $\boldsymbol{T}$ and $\boldsymbol{P}$, and then lower bound
$\mathbb{E}\left[\mathds{1}_{\{T_{i}>0\}}\frac{1}{T_{i}}\right]=\mathbb{E}\left[\frac{1}{T_{i}}\Bigg{|}T_{i}>0\right]\,\mathbb{P}(T_{i}>0)\geq\frac{1}{\mathbb{E}\left[T_{i}\mid T_{i}>0\right]}\,\mathbb{P}(T_{i}>0).$
Finally, note that
$\mathbb{E}[T_{i}\mid T_{i}>0]=\frac{\mathbb{E}T_{i}}{\mathbb{P}(T_{i}>0)}=\frac{m}{n}\frac{1}{\mathbb{P}(T_{i}>0)},$
and $\mathbb{P}(T_{i}>0)=1-\mathbb{P}(T_{i}=0)\geq 1-\exp\left(-\frac{m}{n}\right)$, which concludes the calculation for (I).
Putting everything together,
$\mathbb{E}W_{1}(\widehat{\mu}_{n}(m),\mu_{n})$ $\geq\frac{1}{2}\left[\frac{n}{m}\left(1-\exp\left(-\frac{m}{n}\right)\right)^{2}\left(1-\mathbb{E}\|P\|_{2}^{2}\right)-\exp\left(-\frac{m}{n}\right)\right]$
$\geq\frac{1-\mathbb{E}\|P\|_{2}^{2}}{4}\frac{n}{m},$
provided $m\geq 2n\log\frac{4}{1-\mathbb{E}\|P\|_{2}^{2}}$.
4.3. Proof of Theorem 2.6
The final goal is to obtain an optimal number of cells $n$ given the $m$ reads, which determines the number of reads per cell, in order to minimize $W_{p}(\widehat{\mu}_{n}(m),\mu)$ for $p\in[1,2]$. We only focus on the optimal order of $m$ in terms of $n$ and ignore constants. Observe that a triangle inequality for $W_{p}$ shows
$W_{p}(\widehat{\mu}_{n}(m),\mu_{n})+W_{p}(\mu_{n},\mu).$
Since $W_{p}(\widehat{\mu}_{n}(m),\mu_{n})$ is controlled by Theorem 2.4, let’s pay attention to $W_{p}(\mu_{n},\mu)$.
There have been many papers studying to deduce the tight rate of the convergence of empirical distributions and concentration inequalities of $W_{p}$ for $p\geq 1$. Regarding the state-of-art results in this area, we refers readers to [dudley69, MR2280433, MR2861675, MR3189084, NF_AG_rate_Wasserstein, Niles-Weed_Bach, MR4359822, merigot2021non, MR4255123, MR4441130] and references therein. In the present paper we adopt [Niles-Weed_Bach] since $\Delta_{(d-1)}$ is bounded.
Lemma 4.8.
[Niles-Weed_Bach, Theorem 1, Proposition 20] Let $\mu$ has a support with diameter $D$, and $d^{*}_{p}(\mu)$ be the upper Wasserstein dimension. Let $1\leq p<\infty$. If $k>\max\{d^{*}_{p}(\mu),4\}$, then
$\mathbb{E}[W_{p}(\mu,\mu_{n})]\leq O\left(n^{-\frac{1}{k}}\right).$ (4.10)
Furthermore, for any $t>0$,
$\mathbb{P}\left(W_{p}^{p}(\mu,\mu_{n})\geq t+\mathbb{E}[W_{p}^{p}(\mu,\mu_{n})]\right)\\ \leq\exp\left(-\frac{2nt^{2}}{D^{2p}}\right).$ (4.11)
Remark 4.9.
$k$ is the intrinsic dimension of $\mu$, or the dimension of its support. It is obvious that $k\leq d$. (4.10) implies that if $\mu$ enjoys the intrinsic lower dimensionality, one could expect the faster rate of the convergence of Wasserstein distance of the empirical distributions.
Proof of Theorem 2.6.
Fix $\alpha\in(0,1)$ and $p\in[1,2]$. Assume that there is some $k>0$ such that $\max\{d^{*}_{p}(\mu),4\}<k\leq d$. Since $\text{diam}(\Delta_{(d-1)})=2$, by (4.11), $W_{p}(\mu_{n},\mu)\leq\mathbb{E}W_{p}(\mu_{n},\mu)+n^{-q}$ with probability $1-\exp\left(-\frac{1}{2^{2p-1}}n^{1-2q}\right)$ for any $0<q<\frac{1}{2}$. Let $q=\frac{1}{4}$ for simplicity. Then, with probability $1-\exp\left(-\frac{\sqrt{n}}{2^{2p-1}}\right)$
$W_{p}(\mu_{n},\mu)\leq O(n^{-\frac{1}{k}}).$
Let $\varepsilon=n^{-t}$ for some $t>0$. Given $m$, let $n$ satisfy (2.4). Ignoring the lower order term, a maximum such $n$ should satisfy
$n\geq\left(\frac{c_{*}m}{8(1+\alpha)\mathbb{E}\Arrowvert P\Arrowvert_{0}}\right)^{\frac{1}{1+2t}}.$ (4.12)
Combining (2.5) with (4.11), a triangle inequality implies
$W_{p}(\widehat{\mu}_{n}(m),\mu)\leq O(n^{-t})+O(n^{-\frac{1}{k}})$
with probability $1-O\left(n^{-\frac{\alpha}{2}}\right)=\left(1-O\left(n^{-\frac{\alpha}{2}}\right)\right)\left(1-\exp\left(-\frac{\sqrt{n}}{2^{2p-1}}\right)\right)$. Let’s optimize $O(n^{-t})+O(n^{-\frac{1}{k}})$ by choosing $t$. Let $n(t)$ satisfy (4.12) given $m$ and $t$. Then,
$O(n(t)^{-t})+O(n(t)^{-\frac{1}{k}})$ $=O\left(e^{-t\log n(t)}\right)+O\left(e^{-\frac{1}{k}\log n(t)}\right)=O\left(e^{-\min\{t,\frac{1}{k}\}\log n(t)}\right).$
If $t>\frac{1}{k}$, the order is $O\left(n(t)^{-\frac{1}{k}}\right)$. Notice that $n(t)$ is decreasing with respect to $t$: it is trivial due to (4.12). Hence, $n(t)^{-\frac{1}{k}}$ decreases as $t$ decreases until $t=\frac{1}{k}$. If $t<\frac{1}{k}$, the order is
$O\left(n(t)^{-t}\right)=O\left(e^{-t\log n(t)}\right)=O\left(\exp\left(-\frac{t\left(\log(c_{*}m)-\log\left(8(1+\alpha)\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)\right)}{1+2t}\right)\right).$
It is easy to check that
$\partial_{t}\left(\frac{t\left(\log(c_{*}m)-\log\left(8(1+\alpha)\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)\right)}{1+2t}\right)$
$=\frac{\left(\log(c_{*}m)-\log\left(8(1+\alpha)\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)\right)-2t\left(\log(c_{*}m)-\log\left(8(1+\alpha)\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)\right)}{(1+2t)^{2}}$
$\geq 0$
provided that $\log(c_{*}m)\geq\log\left(8(1+\alpha)\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)$ and $t\leq\frac{1}{2}$. Let $m_{0}(c_{*},\alpha,\mathbb{E}\Arrowvert P\Arrowvert_{0})$ be a positive constant satisfying $\log(c_{*}m)>\log\left(8(1+\alpha)\mathbb{E}\Arrowvert P\Arrowvert_{0}\right)$. Since $m\geq m_{0}$ and $k>4$, increasing $t$ up to $\frac{1}{k}$ decreases $O\left(n(t)^{-t}\right)$. Combining the case for $t>\frac{1}{k}$, we can conclude that the optimal $t=\frac{1}{k}$. ∎
5. Conclusion and future works
In this paper we analyze the distance between an empirical distribution of cells, sampled with scRNA-seq, and the true population. When the total number of reads is fixed, sampling more cells does not always make the empirical distribution $\hat{\mu}$ closer to the true distribution $\mu$. In order to minimize the Wasserstein distance between $\hat{\mu}$ and $\mu$, the number of cells $n$ should scale with the number of reads $m$ like
$n\sim m^{1-\frac{2}{2+k}},$
where $k$ is the intrinsic dimension of the true distribution.
In future work, it would be interesting to analyze optimal sequencing depth for different estimation problems. Fewer reads per cell may be optimal for structured estimation problems such as clustering or shape-constrained density estimation. And adding additional dimensions such as space or time would present interesting analytical opportunities for spatial transcriptomics or trajectory inference. Our results for general empirical distributions should provide a good starting point to analyze these diverse contexts and extract as much information as possible from a limited sequencing budget.
We note that it would be feasible to explore different mathematical models of RNA sequencing, beyond multinomial sampling of reads. In fact, the multinomial distribution plays no crucial role in our analysis: its concentration inequality, (4.2), is the only property used in the proof. For example, one could consider additive or multiplicative noise in addition to multinomial sampling. This could model amplification bias. In our analysis, by contrast, each gene expression profile $P_{i}$ is a pure gene expression driven by $\mu$, and during sequencing $P_{i}$ is kept clean. However, possibly $P_{i}$’s are contaminated by unknown noise. If so, instead of accessing clean $P$, a contaminated gene expression, $Q$, is only accessible to sequencing. Mathematically, it is given as $Q=P+Z$ where $Z$ is an independent random vector. For this case, can we also give the same conclusion? We believe that to deduce a sufficient condition to guarantee the non-trivial detection of the true signal $P$ from $Q$ by sequencing is one of promising future works for understanding many important biological questions.
Finally, one could ask the same question under high-dimensional setting. In our model, the ambient and intrinsic dimensions $d$ and $k$ are fixed, and only $n$ and $m$ grow. In various high-dimensional statistics, however, $d$ also grows as $n$ does (also $k$ sometimes grows but slower than $d$). What happens $d$ also grows? In this case, $\frac{m}{n}\to\alpha$ and $\frac{d}{n}\to\beta$. Under this setting, can the Wasserstein distance between $\widehat{\mu}_{n}(m)$ and $\mu$ vanish? Is there any phase transition of the Wasserstein distance in terms of $\alpha$ and $\beta$? We believe it is a new type of high-dimensional problem of interest both theoretically and practically.
Acknowledgement
JK thanks to PIMS postdoctoral fellowship through the Kantorovich Initiative PIMS Research Network (PRN) as well as National Science Foundation grant NSF-DMS 2133244. GS and SK are supported by a CIHR PG and an NSERC DG. We thank PIMS for their generous support; report identifier PIMS-20240921-PRN01.
[…]
Appendix A Experiment details
In this section, we provide further details on the dataset, preprocessing steps, and software packages used in the experiments presented in Section 3.
A.1. Dataset and preprocessing
The reprogramming dataset is published in [schiebingerOptimalTransportAnalysisSingleCell2019], and is publicly available through the NCBI Gene Expression Omnibus with identification number GSE122662. We use the counts matrices from the files GSM3195748_D15.5_serum_C1_gene_bc_mat.h5 and
GSM3195749_D15.5_serum_C2_gene_bc_mat.h5. Data processing is done in python, using the packages Scanpy [wolf2018scanpy] and Anndata [anndata]. Scanpy’s function read_10x_h5 is used to read in the data files, and the resulting Anndata objects are concatenated via an inner join. The result is a (annotated) cells-by-genes counts matrix of unique molecular identifier (UMI) reads.
We preprocess this data by first filtering out genes that are expressed in fewer than 10 cells, and then subsetting to the top $d=1000$ highly variable genes (as determined by Scanpy’s highly_variable_genes function with flavor = "seurat_v3"). The resulting preprocessed counts matrix $\boldsymbol{C}$ is then used as described in Section 3. Note that we do not use any nonlinear transformation on the data, such as the standard log1p transform.
A.2. Describing the constructed population distribution $\mu$
Recall that $\mu$ is constructed as
$\mu=\frac{1}{N}\sum_{\ell=1}^{N}\delta_{C_{\ell}/\|C_{\ell}\|_{1}}$ (A.1)
where $C_{\ell}$ is the $\ell$-th row of the (preprocessed) counts matrix $\boldsymbol{C}$. As we keep only the top 1000 highly variable genes, the ambient dimension is $d=1000$. Furthermore, as already stated in Section 3, $N=4851$ and $\mathbb{E}_{P\sim\mu}\left\|{P}\right\|_{0}\simeq 175$. Finally, the intrinsic dimension of $\mu$, approximated using the PCA heuristic (Figure A.1), is $k=83$.
Figure A.1. The first 83 principal components of $\mu$ capture 95% of its variance. Here, $\mu$ is computed from the reprogramming dataset.
A.3. Constructing lower-dimensional distributions $\mu_{k}$
Recall the lower-dimensional synthetic distributions $\mu_{k}$ from Section 3.3. Here, we describe our procedure for constructing those distributions from the reprogramming dataset.
The support of $\mu$ as defined in (A.1) is a finite set of vectors in the simplex; it can be packaged as a matrix $M\in\mathbb{R}_{+}^{N\times d}$ whose rows are the atoms of $\mu$. Now, consider a rank-$r$ non-negative matrix factorization $M\simeq WH$, where $W\in\mathbb{R}_{+}^{N\times r}$ and $H\in\mathbb{R}_{+}^{r\times d}$. The rank of $M_{r}:=WH$ is at most $r$, and we can rescale the rows of $M_{r}$ to the simplex without changing the rank. The result is a set of vectors in the simplex that is also contained in an $r$-dimensional affine space. Hence, the rows of this rescaled version of $M_{r}$ form the support of our at-most $r$-dimensional synthetic distribution. We compute its intrinsic dimension $k$ via the PCA heuristic again, noting that $k\leq r$ necessarily. Some statistics are reported in Table 1.
Rank $r$ Relative error Intrinsic dimension $k$ Average sparsity $\mathbb{E}_{P\sim\mu_{k}}\|P\|_{0}$
10 0.45 7 987
30 0.30 15 976
40 0.27 20 981
50 0.24 25 981
Table 1. Statistics of the synthetic distributions $\mu_{k}$. The rank $r$ is the one used for matrix factorization, and the relative error of the factorization is defined as $\left\|{\bar{M_{r}}-M}\right\|_{F}/\left\|{M}\right\|_{F}$ where $\bar{M_{r}}$ is the rescaled version of $WH$. The intrinsic dimension $k$ is computed via the PCA heuristic. Finally, recall that $d=1000$ is the upper bound on the average sparsities.
A.4. Computing optimal transport
The experiments of Section 3 involve computing Wasserstein distances between finitely supported atomic measures. The Python Optimal Transport package [flamary2021POT] provides various tools for this. We used the function emd2, which involves no additional regularization.